OpenStudy (anonymous):

Suppose that f(x+h)-f(x)/h=-2h(x+5)-h^2/h(x+h+5)^2(x+5)^2 Find the slope m of the tangent line at x=4 . m=

OpenStudy (anonymous):

The slope of the tangent line is the derivative function with x=4. Are you familiar with derivatives or new to them? Have you learned the shorter version of the power rule for derivatives or are you still working with this? Limit definition of a derivative: \[ \huge \lim_{{\triangle x} \rightarrow 0} \frac{f(x+\triangle x)+f(x)}{\triangle x} \]

OpenStudy (anonymous):

h = \( \triangle \)x for your problem, same thing, different name

OpenStudy (anonymous):

i am still trying to figure out derivatives

OpenStudy (anonymous):

And the second question I asked? o_O Class progress = ?

OpenStudy (anonymous):

we have yet to get to it but i was working ahead

OpenStudy (anonymous):

I could teach you the shortcuts, but that might mess you up on an exam depending on the teacher you have, if they want to do it the long way until THEY teach you the shortcuts

OpenStudy (anonymous):

sounds good

OpenStudy (anonymous):

Alrighty, let's rewrite that bad boy first, you're going to have to show parenthesis a bit more clearly because currently you have (which is not equals): \[ \frac{f(x+h)-f(x)}{h} \neq-2h(x+5)-h^2/h(x+h+5)^2(x+5)^2 \]

OpenStudy (anonymous):

What's on the left is golden, but whatever you have on the right I need you to rewrite please. Carefully use parenthesis to indicate order of operation the way you meant it. Cool?

OpenStudy (anonymous):

i will note that thanks.

OpenStudy (anonymous):

can u see the file i attached?

OpenStudy (anonymous):

The trick is basically this, you have some function getting x input plus a little bit more and the difference of that with the little bit more Remember this? \[slope=\frac{rise}{run}=\frac{\triangle y}{\triangle x}\] h = change in x = \(\triangle\)x

OpenStudy (anonymous):

I see it. One moment...

OpenStudy (anonymous):

Do you remember talking about average slope?

OpenStudy (anonymous):

\[m_{average}=\frac{\triangle y}{\triangle x}=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

Well what you're doing is making the differences in the two x coordinates get very very small, right down to nothing at all (limit as \(\triangle\)x goes to zero) so you get the exact INSTANTANEOUS slope, instead of average. Which is that very definition of the derivative. With the right hand side of the equation shown in your problem, have you tried to simplify it yet? If not, you need to do that first. You're trying to make stuff cancel out, the h's specifically